# Modeling of Proton Interaction with Organic Polymers: Implications for Cancer Therapy and Beyond

### 05 Jan 2024 v1

@@ INTRODUCTION
Radiation transport simulation codes such as FLUKA, MCNP and TOPAS are widely used for dose estimates in radiation therapy planning [23, 24, 25]. The simulation result enables dose calculations in the planning process, and its quality reflects the quality of fundamental data tables, such as stopping power, scattering cross sections, ionization probabilities, energy loss, and lateral straggling, among others.
@@
In this framework, knowledge of the energy-loss function (ELF) of materials is essential. The Penn approach [45] introduced an algorithm to determine the electron inelastic mean free paths (IMFP) by utilizing a model dielectric function derived from the experimental ELF specific to the material under investigation. The same model has been applied to estimate the electron stopping power in various materials [46] and has been extended to calculate the non-linear stopping power of ions [44]. This extension involves using the ELF to appropriately weight contributions from different electron gas components within a statistical ensemble that characterizes the material of interest. The ELF functions at the optical limit can be found for different materials elsewhere [47].
@@ METHODS
The positive background density of the jellium with radius $R_{\rm cl}$ is defined by $n_{0}^{+}({\bf r})=n_{0}^{+}\Theta(r-R_{\rm cl})$, where $\Theta(x)$ denotes the Heaviside function. The electronic density of the cluster is determined by the Wigner-Seitz radius, $r_{s}$ ($4\pi r_{s}^{3}/3=1/n_{0}$). The total number of electrons in the cluster, $N_{e}$, is then given by $N_{e}=(R_{cl}/r_{s})^{3}$.
@@
Figure 2: Comparative electronic stopping cross-section (SCS) of protons within PE polymer. The blue short dash-dot line is the TDDFT results using a FEG unique ($r_{s}=1.75$ au). The cyan short dash line is the TDDFT-Penn approach result. Experimental data (uppercase letters) are within the Bragg peak region [66]. Semi-empirical models ICRU49 [63] and SRIM-2013 [65] are depicted.
@@
Figure 3: Illustration of electronic stopping cross-section (SCS) for protons in PS polymer. TDDFT results employing a FEG unique ($r_{s}=1.66$ au) are represented by the blue short dash-dot line. The cyan short dash line is the result TDDFT-Penn approach. Experimental data (uppercase letters) are concentrated in the Bragg peak area. Dielectric formalism results are indicated by the purple dash-dot line [62]. Semi-empirical models ICRU49 [63] and SRIM-2013 [65] are showcased.
@@
Notably, the TDDFT-Penn results for PS (refer to Figure 3) exhibit a better agreement with the experimental data compared to SRIM-2013. SRIM-2013 employs Bragg’s rule, resulting in an excitation energy (I) for PS of 65.5 eV [68]. However, electron energy loss spectroscopy (EELS) data from experiments [61] for the compound PS suggest a lower I value (59.3 eV). This observation explains the higher values obtained with our approach in the Bragg peak region.
SCS results for P2VP and PA (refer to Figures 4 and 5) exhibit good agreement with SRIM-2013.
Figure 4: Display of electronic stopping cross-section (SCS) in P2VP polymer for protons. The blue short dash-dot line represents TDDFT results utilizing a unique FEG ($r_{s}=1.66$ au). The cyan short dash line is the result TDDFT-Penn approach. The purple dash-dot line signifies results based on dielectric formalism [62]. Semi-empirical model SRIM-2013 [65] is included.
However, it is important to note that we did not find any experimental data for comparison in this context. Additionally, the dielectric approach, employing a similar
Figure 5: Depiction of electronic stopping cross-section (SCS) of protons interacting with PA polymer. TDDFT results using a FEG unique ($r_{s}=1.62$ au) are represented by the blue short dash-dot line. The cyan short dash line is the result TDDFT-Penn approach. Results based on dielectric formalism are represented by the purple dash-dot line [62]. Semi-empirical model SRIM-2013 [65] is presented.
ELF function as input [62], deviates significantly below the Bragg peak.
@@
Figure 6: Comparison of electronic stopping cross-section (SCS) for protons interacting with PMMA polymer. TDDFT results using a FEG unique ($r_{s}=1.74$ au) are represented by the short blue dashed line. The cyan short dashed line is the result of the TDDFT-Penn approach. Results based on the dielectric formalism are displayed on the purple dashed-dotted line [62]. Semi-empirical models ICRU49 [63] and SRIM-2013 [65] with Bragg rule (red dashed line) and CAB (dotted green line) correction are showcased.
@@
In an analogous fashion, in TiN compounds there is a transfer of 1.51 electrons from titanium to nitrogen [40], and the transferred charges. The transfer of charges is expected to be more noticeable in the double bonds between carbon and oxygen in polymers like PMMA and PI. Therefore, our results suggest the possibility of charge transfer occurring from carbon to oxygen or nitrogen on these polymers. If this transfer is indeed occurring, employing Bragg’s rule will likely lead to a decreased accuracy in predicting the SCS.
Figure 7: Visualization of electronic stopping cross-section (SCS) for protons interacting with PI polymer. TDDFT results using a FEG unique ($r_{s}=1.61$ au) are represented by the blue short dash-dot line. The cyan short dash line is the result TDDFT-Penn approach. The purple dash-dot line represents results based on dielectric formalism [62]. Semi-empirical models ICRU37 [64] and SRIM-2013 [65] with Bragg rule (red dashed line) and CAB (dotted green line) correction are presented.
Indeed, the SRIM-2013 results (red dashed line) for the PMMA and PI compounds are numerically lower in the Bragg peak region compared to TDDFT-Penn predictions.
Theorists have been working for decades to develop methods that can model the physical processes responsible for electronic SCS. Although these methods have been successful, they have yet to be able to cover a wide range of energy using a single approach. Some models explain a limited energy range, while others achieve good agreement with experimental data by treating inner and valence electrons differently. In this work, we present a theoretical framework based on ab initio calculations of different FEGs and the Penn method. This framework achieves excellent agreement with experimental data and reference data tables across a wide range of energy, using a single approach for all electrons.
@@
To showcase the efficiency of our approach, we conducted research on the stopping power of organic polymers for protons. In general, a single FEG model is not enough to accurately describe the stopping power at the Bragg peak. Our new method provides a theoretical framework to tackle electron density variations. We have demonstrated the effectiveness of our method by obtaining excellent agreement with the experimental data and reference data tables for polymers such as PE, PS, P2VP, PA, PMMA and PI (as shown in Figures 2 to 7). Our findings highlight the importance of considering the intricate electronic structures of polymers in the theoretical modeling of stopping power.
@@
These findings emphasize the importance of considering the complex electronic structures of polymers in predicting stopping power accurately. The agreement between the TDDFT-Penn approach and the experimental data, which are indicated by uppercase letters in Figures 2 and 3, demonstrates the precision of the present theoretical framework. However, further validation is necessary, particularly for polymers with limited data.
@@
Differences in predicted mean excitation energies and SCS magnitudes between the TDDFT-Penn approach and semi-empirical models such as SRIM-2013, particularly in the Bragg peak region, highlight the potential influence of molecular structure on these predictions. The distinct chemical compositions of PMMA and PI, which have varying electronegativities in their bonds, likely contribute to charge transfer effects not adequately accounted for by Bragg’s rule. The primary consequence of the breakdown of Bragg’s rule is a reduction in the mean excitation energy by approximately 10% and a decrease in the projectile range by about 3 mm for 200 MeV protons.
@@
In conclusion, the results of this research challenge established assumptions and emphasize the need for precise modeling of materials with complex electronic structures. This study enhances our understanding of ion-polymer interactions and provides a solid foundation for future applications in proton therapy and other fields where accurate predictions of stopping power are essential.
ACKNOWLEDGMENTS

### 25 Jan 2024 v2

@@ Abstract
In this framework, knowledge of the energy-loss function (ELF) of materials is essential. The Penn approach [42] introduced an algorithm to determine the electron inelastic mean free paths (IMFP) by utilizing a model dielectric function derived from the experimental ELF specific to the material under investigation. The same model has been applied to estimate the electron stopping power in various materials [43] and has been extended to calculate the non-linear stopping power of ions [41]. This extension involves using the ELF to appropriately weight contributions from different electron gas components within a statistical ensemble that characterizes the material of interest. ELF functions at the optical limit can be found for different materials elsewhere [44].
@@
The positive background density of the jellium with radius $R_{\rm cl}$ is defined by $n_{0}^{+}({\bf r})=n_{0}^{+}\Theta(R_{\rm-rcl})$, where $\Theta(x)$ denotes the Heaviside function. The electronic density of the cluster is determined by the Wigner-Seitz radius, $r_{s}$ ($4\pi r_{s}^{3}/3=1/n_{0}$). The total number of electrons in the cluster, $N_{e}$, is then given by $N_{e}=(R_{cl}/r_{s})^{3}$.
@@
Figure 2: Proton SCS in PE polymer. TDDFT results with a unique FEG ($r_{s}=1.75$ au) are shown in the blue dash-dot line, and the TDDFT-Penn is in the cyan short dash line. Experimental data (uppercase letters) at Bragg peak [63]. Semi-empirical models ICRU49 [60] and SRIM-2013 [62] presented.
@@
In particular, the TDDFT-Penn results for PS (refer to Figure 3) agree better with the experimental data compared to SRIM-2013. SRIM-2013 employs Bragg’s rule, resulting in an excitation energy (I) for PS of 65.5 eV [65]. However, electron energy loss spectroscopy (EELS) data from experiments [58] for the compound PS suggest a lower I value (59.3 eV). This observation explains the higher values obtained with our approach in the Bragg peak region.
Figure 3: Proton SCS in PS polymer. TDDFT results using a unique FEG ($r_{s}=1.66$ au) and TDDFT-Penn. Experimental data (uppercase letters) concentrated at Bragg peak. Dielectric formalism results in purple dash-dot line [47]. Semi-empirical models ICRU49 [60] and SRIM-2013 [62] showcased.
SCS results for P2VP and PA (see Figures 4 and 5) agree well with SRIM-2013. However, it is important to note that we did not find any experimental data for comparison in this context. Furthermore, the dielectric approach, employing a similar ELF function as input [59], deviates significantly below the Bragg peak.
Figure 4: Proton SCS in P2VP polymer. TDDFT results with a unique FEG ($r_{s}=1.66$ au) and TDDFT-Penn. Results based on dielectric formalism [59]. Semi-empirical model SRIM-2013 [62] included.
@@
Figure 5: Proton SCS in PA polymer. TDDFT results with a unique FEG ($r_{s}=1.62$ au) and TDDFT-Penn. Results based on dielectric formalism [59]. Semi-empirical model SRIM-2013 [62] presented.
As can be observed in Figures 2 to 7 the differences between TDDFT-Penn and SRIM-2003 are small, but more pronounced for PMMA and PI at the Bragg peak. They can be attributed to a stronger breakdown of Bragg’s rule as a result of the molecular structures of PMMA and PI, both of which feature bonds between C and O, as well as N.
Figure 6: Proton SCS in PMMA polymer. TDDFT results with a unique FEG ($r_{s}=1.74$ au) and TDDFT-Penn. Dielectric formalism results [59]. Semi-empirical models ICRU49 [60] and SRIM-2013 [62] with Bragg’s rule (red dashed line) and CAB (dotted green line) correction showcased.
In contrast to other depicted polymers, these two have particularities in their chemical structures.
Figure 7: Proton SCS in PI polymer. TDDFT results with a unique FEG ($r_{s}=1.61$ au) and TDDFT-Penn. Results based on dielectric formalism [59]. Semi-empirical models ICRU37 [61] and SRIM-2013 [62] with Bragg’s rule and CAB correction presented.
@@
Analogously, in TiN compounds, there is a transfer of 1.51 electrons from titanium to nitrogen [37], and the transferred charges. The transfer of charges is expected to be more noticeable in the double bonds between carbon and oxygen in polymers like PMMA and PI. Therefore, our results suggest the possibility of charge transfer occurring from carbon to oxygen or nitrogen on these polymers. If this transfer is indeed occurring, employing Bragg’s rule will likely lead to a decreased accuracy in predicting the SCS. Indeed, the SRIM-2013 results (red dashed line) for the PMMA and PI compounds are numerically lower in the Bragg peak region compared to TDDFT-Penn predictions.
Decades of theoretical efforts to model electronic SCS processes have led to various methods. Despite successes, current approaches fall short of uniformly covering a wide energy range. Some models focus on specific energy ranges, while others achieve agreement by treating inner and valence electrons differently. Our ab initio framework, using various FEG models and the Penn method, shows excellent agreement across a broad energy spectrum for all electrons.
@@
Exploring proton-stopping power in organic polymers, we highlight the challenge at the Bragg peak for traditional single-FEG models. Our method effectively addresses electron density variations, which shows agreement with experimental data and reference tables for various polymers. The importance of considering intricate electronic structures in theoretical stopping power modeling for polymers is underscored.
@@
The accurate prediction of stopping power requires the inclusion of complex electronic structures. The TDDFT-Penn approach’s agreement with experimental data underscores the precision of our theoretical framework, necessitating further validation, particularly for polymers with limited data.
@@
Discrepancies between the TDDFT-Penn approach and semi-empirical models, such as SRIM-2013, underscore the potential impact of molecular structure on predictions. Varied electronegativities in PMMA and PI bonds contribute to charge transfer effects not fully accounted for by Bragg’s rule. The primary consequence is a 10% reduction in mean excitation energy and a 3 mm decrease in the projectile range for 200 MeV protons.
@@
In conclusion, this research challenges assumptions, emphasizing precise modeling for materials with complex electronic structures. Enhancing understanding of ion-polymer interactions, this study provides a robust foundation for applications in proton therapy and fields requiring accurate stopping power predictions.
Acknowledgements.